# Monotone Sequences of Metric Spaces with Compact Limits

### v1

@@ Abstract.
In this paper, we consider a fixed metric space (possibly an oriented Riemannian manifold with boundary) with an increasing sequence of distance functions and a uniform upper bound on diameter. When the metric space endowed with the pointwise limit of these distances is compact, then there is uniform and Gromov-Hausdorff (GH) convergence to this limit. When the metric space also has an integral current structure of uniformly bounded total mass (as is true for an oriented Riemannian manifold with boundary that has a uniform bound on total volume), we prove volume preserving intrinsic flat convergence to a subset of the GH limit whose closure is the whole GH limit. We provide a review of all notions and have a list of open questions at the end. Dedicated to Xiaochun Rong.
@@ 1. Introduction
Our goal is to teach the notions of Gromov-Hausdorff [GH] and Sormani-Wenger intrinsic flat [SWIF] convergence while proving a new theorem that has no assumptions on curvature. The notion of GH distance between metric spaces was first introduced by Edwards in [Edw75] and deeply explored by Gromov in [Gro07] and [Gro81]. See Rong’s 2010 survey [Ron10] for many applications of GH convergence to sequences of Riemannian manifolds with curvature bounds.
@@
The work in this paper is inspired by a theorem in the appendix of [HLS17] by Huang-Lee-Sormani (which applies to sequences with biLipschitz bounds on their distances) and a theorem within [APS24] by Allen-Perales-Sormani (which assumes only volume converging from above and distance from below but requires the limit space to be a compact smooth oriented Riemannian manifold). Here we will only assume the limit space is a compact metric space but we add the hypothesis that the distances are monotone increasing. Our result will be applied by Sormani-Tian in [ST24] to prove SWIF and GH convergence to the extreme limits constructed by Sormani-Tian-Wang in [STW24].
@@
Given a compact Riemannian manifold, $(M^{m},g_{0})$, possibly with boundary, with a monotone increasing sequence of Riemannian metric tensors $g_{j}$ such that
@@
(1.2) $\operatorname{diam}_{g_{j}}(M)\leq D_{0}.$
@@
If the metric space $(M,d_{\infty})$ is a compact metric space, then $d_{j}\to d_{\infty}$ uniformly and
@@
If, in addition, $M$ is an oriented Riemannian manifold with uniformly bounded total volume,
(1.4) $\operatorname{vol}_{j}(M)\leq V_{0}\textrm{ and }\operatorname{vol}_{j}(\partial M)\leq A_{0}$
then we have volume preserving intrinsic flat convergence
(1.5) $(M,d_{j},[[M]])\stackrel{{\scriptstyle\mathcal{VF}}}{{\longrightarrow}}(M_{\infty},d_{\infty},T_{\infty}).$
where $T_{\infty}$ is an integral current on $(M,d_{\infty})$ such that $T_{\infty}=[[M]]$ viewed as an integral current on $(M,d_{j})$ and
@@
In Section 3 we review the definition of Gromov-Hausdorff convergence (Definition 3.1). We then prove the GH and uniform convergence part of Theorem 1.1. In fact we prove a more general theorem concerning a metric space with a monotone increasing sequence of distance functions that converge pointwise to a compact limit in Theorem 3.2. We show the sequence converges uniformly and in the GH sense. We prove this theorem using different techniques than would have been used by Gromov in order to prepare for the SWIF convergence part of the paper. In particular, we construct a common metric space for the converging sequence that has special properties in Proposition 3.9 and Lemma 3.10. In Remark 3.8 we explain that Example 2.6 does not have a Gromov-Hausdorff limit.
In Section 4 we prove the SWIF convergence part of Theorem 1.1. In fact we prove a more general SWIF convergence theorem concerning the convergence of an integral current space with increasing distance functions [Theorem 4.16]. We begin by reviewing De Giorgi’s notion of a Lipschitz tuple [DG95], the notion of a Lipschitz chart, and Ambrosio-Kirchheim’s notion of an integral current and weak convergence [AK00], proving lemmas about how these objects behave for monotone increasing sequences of distances. We review Sormani-Wenger’s definition of an integral current space, including $(M,d,[[M]])$, and their definition of intrinsic flat distance. We then state Theorem 4.16 and apply it to complete the proof of Theorem 1.1 in Section 4.8. Finally we prove Theorem 4.16, first finding a converging subsequence and a limit current structure in Proposition 4.17 and completing the proof in Section 4.10.
@@
The authors would like to thank Wenchuan Tian and Changliang Wang for inspiring discussions of examples.
@@ 2. Pointwise Convergence of Monotone Increasing Distance Functions
Here we consider metric spaces of finite diameter,
@@ 2.1. Riemannian Distances
Remark 2.4.
As a consequence of Lemma 2.3, a Riemannian manifold with a monotone increasing sequence of metric tensors as in (1.1), has a monotone increasing sequence of distance functions as in Definition 2.2. An easy way to construct such a sequence on the sphere, ${\mathbb{S}}^{m}$, is to take a sequence of graphs over spheres, $g_{j}=g_{{\mathbb{S}}^{m}}+dh_{j}^{2}$ where $g_{{\mathbb{S}}^{m}}$ is the standard round metric on the sphere and $h_{j}:{\mathbb{S}}^{m}\to[0,1]$ has $|\nabla h_{j}|\leq|\nabla h_{j+1}|\textrm{ on }{\mathbb{S}}^{m},$ so that $L_{g_{j}}(C)\leq L_{g_{j+1}}(C)$ for all curves $C$.
@@ 2.3. Example without Uniform Convergence
Even if $(X,d_{j})$ is a monotone increasing sequence of compact metric spaces with a uniform upper bound on diameter pointwise convergence does not imply uniform convergence. See Example 2.6 depicted in Figure 2.1.
@@
Here we adapt an example from Sormani-Wenger [SW11] as in Figure 2.1. Consider a sequence of distinct points $p_{j}\to p_{\infty}$ lying on the equator in the standard round sphere $({\mathbb{S}}^{2},g_{\mathbb{S}^{2}})$. Let $r_{j}>0$ be decreasing radii such that the balls $B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$ are disjoint and do not contain $p_{\infty}$. We define an increasing sequence of metric tensors
@@
as in Remark 2.4 inductively starting with constant $h_{0}=0$ and taking
(2.21) $h_{j+1}(x)$ $=$ $h_{j}(x)\qquad\textrm{ for }x\in{\mathbb{S}}^{2}\setminus B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$
(2.22) $=$ $H(d_{\mathbb{S}^{2}}(p_{j},x)/r_{j})\quad\textrm{ for }x\in B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$
@@
So we have added increasingly thin wells in place of each of the balls. For all $i\leq j$, the $g_{j}$ radial distance from $p_{i}$ to any $q_{i}\in\partial B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$,
@@ 3.2. Uniform Convergence
By Lemma 2.5, we have $d_{j}\leq d_{\infty}$ as functions on $X\times X$. By Lemma 3.5, the maps $d_{j}:X\times X\to[0,D_{0}]$ are bounded $1$-Lipschitz functions on the metric space $(X\times X,d_{sum,\infty,\infty})$, which is compact by Lemma 3.4.
@@
(3.11) $d_{\infty}(x,y)-\epsilon<d_{j_{k}}(x,y)\leq d_{j}(x,y)\leq d_{\infty}(x,y)\qquad\forall x,y\in X\times X.$
@@ 3.3. Constructing a Common Metric Space
3.4. Proof of GH Convergence
Theorem 3.2 is now easy to prove:
@@ 3.4. Proof of GH Convergence
First we apply Prop 3.6 to get uniform convergence from below. Then we apply Proposition 3.7 to $(X,d_{j})$ and $(X,d_{\infty})$ to see that
@@
Note that in the proof of Theorem 3.2 it was essential that we assumed $(X,d_{\infty})$ is compact, as seen in Example 2.6. This example has the monotonicity but no Gromov-Hausdorff limit. In fact Gromov proved that if a sequence of compact metric spaces has a GH limit then the entire sequence can be embedded via distance preserving maps into a common metric space, $(Z,d_{Z})$ [Gro81]. However, the Riemannian manifolds in Example 2.6 have an unbounded number of disjoint balls of radius $1$, so this is impossible.
@@ 3.5. A Common Compact Metric Space for the Whole Sequence
so the inclusion maps
(3.39) $\zeta_{j}:(Z_{j},d_{Z_{j}})\to(Z,d_{Z})$
are distance preserving is a distance function. In addition, taking
@@ 4.1. Lipschitz Functions and Tuples on Metric Spaces
We say that a function, $F:(X,d_{X})\to(Y,d_{Y})$ is $K$-Lipschitz if
(4.1) $d_{Y}(F(x_{1},x_{2}))\leq Kd_{X}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in X$
@@ 4.2. Charts into our Metric Space
This is needed to understand integral structures on $(X,d_{j})$ vs $(X,d_{j+1})$ and $(X,d_{\infty})$.
@@
Sometimes in the definition of a rectifiable space or current it is written that the charts $\psi_{i}:U_{i}\to X$ are bi-Lipschitz. This is true by Lemma 4 in Kirchheim’s [Kir94] as soon as one has a countable collection of Lipschitz charts (cf. Lemma 4.1 of [AK00]). The bi-Lipschitz charts are not necessarily the same collection of charts. So even if the charts $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{\mathbb{R}^{m}})\to(X,d_{b})$ are bi-Lipschitz onto their images, they are only Lipschitz as charts into $(X,d_{a})$ not bi-Lipschitz, and so one would need to apply Kirchheim’s theorem to produce bi-Lipschitz charts into $(X,d_{b})$.
@@ 4.3. Ambrosio-Kirchheim’s integral currents on complete metric spaces
An $m$ dimensional integer rectifiable current on a complete metric space, $(Z,d_{Z})$, is a multilinear functional, $T$, defined on Lipschitz tuples on $(Z,d_{Z})$ that has parametrization which is an atlas of Lipschitz charts
(4.12) $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}}\to(Z,d_{Z})$
@@
The integral is well defined because charts are Lipschitz and thus differentiable almost everywhere. The weighted sum of the integrals defining $T(\pi_{0},...,\pi_{m})$ is finite because the weighted sum of the Hausdorff measures in (4.13), $\max\{\pi_{0}\}$, and the product $\operatorname{Lip}(\pi_{j})$ for $j=1,\ldots,m$ can be used to bound the integrals. Warning: the Lipschitz charts and the Hausdorff measures depend on the distance function on $Z$. See Lemma 4.7 below.
By Stoke’s Theorem, an oriented Riemannian manifold of dimension $m$, $M$, has
@@
So it is natural to define the boundary of a current by
(4.17) $\partial T(\pi_{1},...,\pi_{m})=T(1,\pi_{1},...,\pi_{m}).$
An integral current is an integer rectifiable current whose boundary is also integer rectifiable. Be warned however that the parametrization of the boundary, $\partial T$, is not necessarily found by taking the boundaries of the atlas of charts parametrizing, $T$.
@@
The same can be done for $\partial S$. This also gives us the total mass estimate in (4.26). ∎
Note in particular the above lemma implies that distance preserving maps push forward currents conserving their masses, boundaries, and boundary masses.
@@
The converse of Lemma 4.7 does not hold. If $T$ is an integral current on $(X,d_{a})$, it does not necessarily define an integral current on $(X,d_{b})$ for $d_{b}\geq d_{a}$. It is a multilinear functional acting on tuples in $(X,d_{b})$ because these are also tuples on $(X,d_{a})$ by Lemma 4.1. However it does not necessarily have finite mass. In addition, a parametrization of $T$ by Lipschitz charts into $(X,d_{a})$ might not give Lipschitz charts into $(X,d_{b})$.
@@
If we have an integral current $T$ on $(X,d_{a})$ and there is an integral current, $S$, on $(X,d_{b})$ where $d_{b}\geq d_{a}$ such that
(4.31) $S(\pi_{0},\pi_{1},...,\pi_{m})=T(\pi_{0},\pi_{1},...,\pi_{m}).$
for any tuple of $d_{a}$ Lipschitz functions on $X$, then we say that $S=T$ viewed as an integral current on $(M,d_{a})$. This means we can use any parametrizations of $S$ and $\partial S$ with charts into $(X,d_{b})$ as parametrizations for $T$ and $\partial T$ with the same charts and weights into $(X,d_{a})$ respectively. So we can view $T$ as an integral current on $(X,d_{b})$.

### v2

@@ Abstract.
In this paper, we consider a fixed metric space (possibly an oriented Riemannian manifold with boundary) with an increasing sequence of distance functions and a uniform upper bound on diameter. When the fixed space endowed with the pointwise limit of these distances is compact, then there is uniform and Gromov-Hausdorff (GH) convergence to this space. When the fixed metric space also has an integral current structure of uniformly bounded total mass (as is true for an oriented Riemannian manifold with boundary that has a uniform bound on total volume), we prove volume preserving intrinsic flat convergence to a subset of the GH limit whose closure is the whole GH limit. We provide a review of all notions and have a list of open questions at the end. Dedicated to Xiaochun Rong.
@@ 1. Introduction
Our goal is to teach the notions of Gromov-Hausdorff (GH) and Sormani-Wenger intrinsic flat (SWIF) convergence while proving a new theorem that has no assumptions on curvature. The notion of GH distance between metric spaces was first introduced by Edwards in [Edw75] and deeply explored by Gromov in [Gro07] and [Gro81]. See Rong’s 2010 survey [Ron10] for many applications of GH convergence to sequences of Riemannian manifolds with curvature bounds.
@@
The work in this paper is inspired by a theorem in the appendix of [HLS17] by Huang-Lee-Sormani (which applies to sequences with biLipschitz bounds on their distances) and a theorem within [APS24] by Allen-Perales-Sormani (which assumes only volume converging from above and distance from below but requires the limit space to be a compact smooth oriented Riemannian manifold). Here we will only assume the limit space is a compact metric space but we add the hypothesis that the distances are monotone increasing. Our result will be applied by Sormani-Tian-Yeung in [STY25] to prove SWIF and GH convergence to the extreme limits constructed by Sormani-Tian-Wang in [STW24].
@@
Given a compact connected Riemannian manifold, $(M^{m},g_{0})$, possibly with boundary, with a monotone increasing sequence of Riemannian metric tensors $g_{j}$ such that
@@
(1.2) $\operatorname{diam}_{g_{j}}(M)\leq D_{0}\qquad\forall j\in{\mathbb{N}}.$
@@
If the metric space $(M,d_{\infty})$ is a compact metric space, then $d_{j}\to d_{\infty}$ uniformly and we have Gromov-Hausdorff (GH) convergence,
@@
If, in addition, $M$ is an oriented manifold with uniform bounds on volume and boundary volume,
(1.4) $\operatorname{vol}_{j}(M)\leq V_{0}\textrm{ and }\operatorname{vol}_{j}(\partial M)\leq A_{0}\qquad\forall j\in{\mathbb{N}}$
then we have volume preserving intrinsic flat ($\mathcal{VF}$) convergence
(1.5) $(M,d_{j},[[M]])\stackrel{{\scriptstyle\mathcal{VF}}}{{\longrightarrow}}(M_{\infty},d_{\infty},T_{\infty}),$
where $T_{\infty}$ is an integral current on $(M,d_{\infty})$ such that “$T_{\infty}=[[M]]$ viewed as an integral current on $(M,d_{j})$” and
@@
In fact, we prove more general statements that do not require $M$ to be a manifold. See Theorem 3.2 within for GH convergence of a monotone sequence of metric spaces with a uniform upper bound on diameter whose distance functions converge pointwise to the distance function of a compact metric space. See Theorem 4.16 and Definition 4.9 within for ${\mathcal{VF}}$ and SWIF convergence. To keep the introduction simple, we do not state these results here.
@@
In Section 3 we review the definition of Gromov-Hausdorff convergence (Definition 3.1). We then prove the GH and uniform convergence part of Theorem 1.1. In fact, we prove a more general theorem concerning a metric space with a monotone increasing sequence of distance functions that converge pointwise to a compact limit in Theorem 3.2. We show that the sequence converges uniformly and in the GH sense. We prove this theorem using different techniques than would have been used by Gromov in order to prepare for the SWIF convergence part of the paper. In particular, we construct a common metric space for the converging sequence that has special properties in Proposition 3.9 and Lemma 3.10. In Remark 3.8 we explain that Example 2.6 does not have a Gromov-Hausdorff limit.
In Section 4 we prove the SWIF convergence part of Theorem 1.1. In fact, we prove a more general SWIF convergence theorem concerning the convergence of an integral current space with increasing distance functions [Theorem 4.16]. We begin by reviewing De Giorgi’s notion of a Lipschitz tuple [DG95], the notion of a Lipschitz chart, and Ambrosio-Kirchheim’s notion of an integral current and weak convergence [AK00], proving lemmas about how these objects behave for monotone increasing sequences of distances. We review Sormani-Wenger’s definition of an integral current space, including $(M,d,[[M]])$, and their definition of intrinsic flat distance. We then state Theorem 4.16 and apply it to complete the proof of Theorem 1.1 in Section 4.8. Finally we prove Theorem 4.16, first finding a converging subsequence and a limit current structure in Proposition 4.17 and completing the proof in Section 4.10 applying Proposition 4.18 to estimate the SWIF distance.
@@
The authors would like to thank Wenchuan Tian and Changliang Wang for inspiring discussions of examples that motivated this paper. We would also like to thank the referee and students, Wai Ho Yeung and Kevin Bui, for careful reading of various parts of the first version of this paper and helpful feedback.
@@ 2. Pointwise Convergence of Monotone Increasing Distance Functions
Here we consider metric spaces $(X,d)$ of finite diameter,
@@ 2.1. Riemannian Distances
Lemma 2.4.
Consider a fixed Riemannian manifold, $(M,g_{0})$, with a sequence of functions, $h_{j}:M\to[0,\infty)$, such that the gradients have magnitudes satisfying,
(2.11) $|\nabla h_{j}|_{g_{0}}\leq|\nabla h_{j+1}|_{g_{0}}$
everywhere on $M$, and have the same direction,
(2.12) $\frac{\nabla h_{j}}{\,\,\,\,|\nabla h_{j}|_{g_{0}}}=\frac{\nabla h_{j+1}}{\,\,\,\,|\nabla h_{j+1}|_{g_{0}}}$
whenever $|\nabla h_{j}|_{g_{0}}>0$. Then the sequence of Riemannian metrics
(2.13) $g_{j}=g_{0}+dh_{j}^{2}$
induced by the graphs of the functions, $h_{j}$, is a monotone increasing sequence of metric tensors, $g_{j}$ as in as in (1.1), and defines a monotone increasing sequence of distance functions as in Definition 2.2.
Proof.
By Lemma 2.3, we need only show (1.1). Recall that
(2.14) $g_{j}(V,V)=g_{0}(V,V)+dh_{j}^{2}(V,V)=g_{0}(V,V)+g_{0}(\nabla h_{j},V)^{2}.$
At points where $|\nabla h_{j}|_{g_{0}}=0$, we thus have
(2.15) $g_{j}(V,V)=g_{0}(V,V)+0\leq g_{0}(V,V)+dh_{j+1}^{2}(V,V)=g_{j+1}(V,V).$
At points where $|\nabla h_{j}|_{g_{0}}>0$, the gradients have the same direction so
(2.16) $g_{j}(V,V)$ $=$ $g_{0}(V,V)+g_{0}(\nabla h_{j},V)^{2}$
(2.17) $=$ $g_{0}(V,V)+g_{0}(\tfrac{\nabla h_{j}}{|\nabla h_{j}|},V)^{2}\,|\nabla h_{j}|^{2}$
(2.18) $=$ $g_{0}(V,V)+g_{0}(\tfrac{\nabla h_{j+1}}{|\nabla h_{j+1}|},V)^{2}\,|\nabla h_{j}|^{2}$
(2.19) $\leq$ $g_{0}(V,V)+g_{0}(\tfrac{\nabla h_{j+1}}{|\nabla h_{j+1}|},V)^{2}\,|\nabla h_{j+1}|^{2}$
(2.20) $=$ $g_{0}(V,V)+dh_{j+1}^{2}(V,V)=g_{j+1}(V,V).$
This completes the proof that (1.1) holds at all points in $M$. ∎
@@ 2.3. Example without Uniform Convergence
Even if $(X,d_{j})$ is a monotone increasing sequence of compact metric spaces with a uniform upper bound on diameter, pointwise convergence does not imply uniform convergence. See Example 2.6 depicted in Figure 2.1.
@@
Here we adapt an example from Sormani-Wenger [SW11] as in Figure 2.1. Consider a sequence of distinct points $p_{j}\to p_{\infty}$ lying on the equator in the standard round sphere $({\mathbb{S}}^{2},g_{\mathbb{S}^{2}})$. Let $r_{j}>0$ be decreasing radii such that the balls $B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$ are disjoint and do not contain $p_{\infty}$. We define a sequence of metric tensors
@@
inductively starting with constant $h_{0}=0$ and taking
(2.31) $h_{j+1}(x)=\begin{cases}h_{j}(x)&\text{if }x\in{\mathbb{S}}^{2}\setminus B_{g_{\mathbb{S}^{2}}}(p_{j+1},r_{j+1})\\ H(d_{\mathbb{S}^{2}}(p_{j+1},x)/r_{j+1})&\text{if }x\in B_{g_{\mathbb{S}^{2}}}(p_{j+1},r_{j+1})\end{cases}$
@@
This implies that, for $x$ outside of the disjoint balls about $p_{0}$ to $p_{i}$, we have
(2.32) $h_{i+1}(x)=h_{i}(x)=...=h_{0}(x)=0.$
Since the balls are disjoint, for all $x\in B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$, $x$ is not in the previous balls, so $h_{j}(x)=0$, so $\nabla h_{j}=0$. Elsewhere, by (2.31) we have $h_{j}=h_{j+1}$ so $\nabla h_{j}=\nabla h_{j+1}$. Thus we can apply Lemma 2.4 to see that we have a monotone increasing sequence of metric tensors.
Note that we have added increasingly thin wells in place of each of the balls. For all $i\leq j$, the $g_{j}$ radial distance from $p_{i}$ to any $q_{i}\in\partial B_{g_{\mathbb{S}^{2}}}(p_{j},r_{j})$,
@@ 3.2. Uniform Convergence
By Lemma 2.5, we have $d_{j}\leq d_{\infty}$ as functions on $X\times X$. By Lemma 3.5, the maps $d_{j}:X\times X\to[0,D_{0}]$ are bounded $1$-Lipschitz functions on the metric space $(X\times X,d_{sum,\infty,\infty})$, which is compact by hypothesis and Lemma 3.4.
@@
(3.11) $d_{\infty}(x,y)-\epsilon<d_{j_{k}}(x,y)\leq d_{j}(x,y)\leq d_{\infty}(x,y)\qquad\forall x,y\in X\times X,$
showing that the whole sequence $d_{j}$ uniformly converges to $d_{\infty}$. ∎
@@ 3.3. Constructing a Common Metric Space
3.4. Proof of Theorem 3.2
GH Convergence is now easy to prove:
@@
First we apply Proposition 3.6 to get uniform convergence of $d_{j}$ to $d_{\infty}$ from below. Then we apply Proposition 3.7 to $(X,d_{j})$ and $(X,d_{\infty})$ to see that
@@
Note that in the proof of Theorem 3.2 it was essential that we assumed $(X,d_{\infty})$ is compact, as seen in Example 2.6. This example has the monotonicity but no Gromov-Hausdorff limit. In fact, Gromov proved that if a sequence of compact metric spaces has a GH limit then the entire sequence can be embedded via distance preserving maps into a common metric space, $(Z,d_{Z})$ [Gro81]. However, the Riemannian manifolds in Example 2.6 have an unbounded number of disjoint balls of radius $1$, so this is impossible.
@@ 3.5. A Common Compact Metric Space for the Whole Sequence
so the inclusion maps,
(3.39) $\zeta_{j}:(Z_{j},d_{Z_{j}})\to(Z,d_{Z}),$
are distance preserving. In addition, taking
@@ 4.1. Lipschitz Functions and Tuples on Metric Spaces
We say that a function, $F:(X,d_{X})\to(Y,d_{Y})$ is $K$-Lipschitz if there exists $K\geq 0$ such that
(4.1) $d_{Y}(F(x_{1},x_{2}))\leq Kd_{X}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in X.$
@@ 4.2. Charts into our Metric Space
We say that a map $\psi:U\subset{\mathbb{R}}^{m}\to X$ from a Borel subset $U$ in ${\mathbb{R}}^{m}$ into a metric space, $(X,d_{X})$, is a Lipschitz chart if there exists $K_{\psi}>0$ such that
(4.8) $d_{X}(\psi(a_{1}),\psi(a_{2}))\leq K_{\psi}|a_{1}-a_{2}|\quad\forall a_{1},a_{2}\subset U.$
These charts will be applied in the next section to define rectifiable and integral currents, and then integral current spaces in the following section.
@@
Sometimes in the definition of a rectifiable space or current it is written that the charts $\psi_{i}:U_{i}\to X$ are bi-Lipschitz. However, once one has a countable collection of Lipschitz charts, one can replace them with a collection of bi-Lipschitz charts by Lemma 4 in Kirchheim’s [Kir94] (cf. Lemma 4.1 of [AK00]). The bi-Lipschitz charts are not necessarily the same collection of charts. If we apply our Lemma 4.5 to a collection of charts $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{\mathbb{R}^{m}})\to(X,d_{b})$ that are bi-Lipschitz onto their images, we conclude they are Lipschitz as charts into $(X,d_{a})$ but cannot determine if they are bi-Lipschitz. We would need to apply Kirchheim’s theorem to produce a new collection of bi-Lipschitz charts into $(X,d_{b})$ if we need them.
@@ 4.3. Ambrosio-Kirchheim’s integral currents on complete metric spaces
An $m$ dimensional integer rectifiable current on a complete metric space, $(Z,d_{Z})$, is a multilinear functional, $T$, defined on Lipschitz tuples on $(Z,d_{Z})$ that has a parametrization which is a countable collection of Lipschitz charts with non-zero Lipschitz constants
(4.13) $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}})\to(Z,d_{Z})$
@@
This integral is well defined because the charts are Lipschitz and thus differentiable almost everywhere. The weighted sum of the integrals defining $T(\pi_{0},...,\pi_{m})$ is finite because the weighted sum of the Hausdorff measures in (4.14), $\max\{\pi_{0}\}$, and the product $\operatorname{Lip}(\pi_{j})$ for $j=1,\ldots,m$ can be used to bound the integrals. Warning: the Lipschitz charts and the Hausdorff measures depend on the distance function on $Z$. See Lemma 4.7 below.
Note that if $(M^{m},g)$ is a compact oriented Riemannian manifold with boundary, with a smooth atlas of oriented charts $\psi_{i}:(W_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}})\to(M^{m},d_{g})$ then we can choose $U_{i}\subset W_{i}$ such that $\psi_{i}(U_{i})$ are disjoint and their union covers $M^{m}$. This defines a natural current structure for $M$ with weight $1$, which agrees with integration of differential $m$ forms:
@@
By Stoke’s Theorem, for any collection of smooth functions, $\pi_{i}:M\to{\mathbb{R}}$, for $i=1,...,m-1$, we have
[…]
To be consistent with Stoke’s Theorem, Ambrosio-Kirchheim define the boundary of a current, $T$, as follows
(4.19) $\partial T(\pi_{1},...,\pi_{m-1})=T(1,\pi_{1},...,\pi_{m-1}),$
for any Lipschitz tuple $(\pi_{1},...,\pi_{m-1})$ on $Z$.
They define an integral current to be an integer rectifiable current whose boundary is also integer rectifiable. Be warned however that the parametrization of the boundary, $\partial T$, is not necessarily found by taking the boundaries of the atlas of charts parametrizing, $T$.
@@
The same holds for $\partial S$ which implies the total mass estimate in (4.28). ∎
Note the above lemma implies that distance preserving maps push forward currents conserving their masses, boundaries, and boundary masses.
@@
The converse of Lemma 4.7 does not hold in general. If $T$ is an integral current on $(X,d_{a})$, it does not necessarily define an integral current on $(X,d_{b})$ for $d_{b}\geq d_{a}$. A parametrization of $T$ by Lipschitz charts into $(X,d_{a})$ might not give Lipschitz charts into $(X,d_{b})$ because that would require a converse to Lemma 4.5 (which we do not have). In addition, $T$ does not necessarily act on tuples in $(X,d_{b})$ because that would require a converse to Lemma 4.1 (which we do not have). Note that if one could prove that either or both of the above were true for some special $T$, one would still need to verify that the mass of $T$ and $\partial T$ are finite with respect to $d_{b}$ to conclude that $T$ is an integral current on $(X,d_{b})$.
@@
If we have an integral current $T$ on $(X,d_{a})$ and there is also an integral current, $S$, on $(X,d_{b})$ where $d_{b}\geq d_{a}$ such that
(4.33) $S(\pi_{0},\pi_{1},...,\pi_{m})=T(\pi_{0},\pi_{1},...,\pi_{m})$
for any tuple of $d_{a}$ Lipschitz functions on $X$, then we say that “$S=T$ viewed as an integral current on $(X,d_{a})$”. That is, we can use any parametrization of $S$ and $\partial S$ with charts into $(X,d_{b})$ as a parametrization for $T$ and $\partial T$ with the same weights and charts into $(X,d_{a})$ respectively.  [^1] Note that Lemma 4.7 applies to $S$ but not $T$ as discussed in Remark 4.8.